# reference.bib Erase, then Redraw: A Novel Data Augmentation Approach for Free Space Detection Using Diffusion Model

### v1

@@ Abstract
Data augmentation is one of the most common tools in deep learning, underpinning many recent advances including tasks such as classification, detection, and semantic segmentation. The standard approach to data augmentation involves simple transformations like rotation and flipping to generate new images. However, these new images often lack diversity along the main semantic dimensions within the data. Traditional data augmentation methods cannot alter high-level semantic attributes such as the presence of vehicles, trees, and buildings in a scene to enhance data diversity. In recent years, the rapid development of generative models has injected new vitality into the field of data augmentation. In this paper, we address the lack of diversity in data augmentation for road detection task by using a pre-trained text-to-image diffusion model to parameterize image-to-image transformations. Our method involves editing images using these diffusion models to change their semantics. In essence, we achieve this goal by erasing instances of real objects from the original dataset and generating new instances with similar semantics in the erased regions using the diffusion model (as shown in Fig. 1), thereby expanding the original dataset. We evaluate our approach on the KITTI road dataset [fritsch2013new] and achieve the best results compared to other data augmentation methods, which demonstrates the effectiveness of our proposed development. Here is our project page: https://sites.google.com/view/data-augmentation.
@@ I INTRODUCTION
In this paper, we propose a novel data augmentation method for free space detection. The method is mainly divided into two steps: first, erasing instance pixels in the background using traditional instance segmentation algorithms such as Mask R-CNN [he2017mask] or a general segmentation algorithm like Segment Anything (SAM) [kirillov2023segment]. Then, the erased regions are locally redrawn using a pre-trained diffusion model to restore the erased parts in the image, as shown in the Fig. 1. During the redrawing process, different linguistic prompts can be used to achieve redrawing of different objects and styles, providing great flexibility. We then test our proposed method on the KITTI road dataset [fritsch2013new], and the experimental results demonstrate the effectiveness of our approach. Our main contributions are as follows:
@@
We conduct comprehensive experiments on the KITTI road dataset, and the results demonstrate that our data augmentation method achieves the best performance in the free space detection task.
We integrate our data augmentation method into a user-friendly GUI interface, equipped with both automatic and manual augmented data generation functions to facilitate easier use for users.
@@ III-B Our Approach
For the “erasing” process, i.e., the masks generation step, we utilized SAM [kirillov2023segment] model, proposed by Menta AI. As a foundational model for image segmentation, SAM demonstrates strong generalization capabilities and performs well across different data domains. We employ two erasure modes in our study. The first mode is the single-point mode on our GUI, where objects to be erased are manually selected. In the second method, we use the approach proposed in [li2023clip], where text prompts specifying the objects to be erased are used to generate candidate points on the original image. These candidate points are then utilized as point prompts for the SAM model to erase the objects in the original image. Both methods are integrated into our GUI interface for the convenience of researchers.
@@
Network | Network Input | Augmentation Method | Accuracy | Precision | Recall | F1-Score | mIoU |
PLARD [chen2019progressive] | RGB + LiDAR | Standard | 95.61 | 92.08 | 91.09 | 91.58 | 85.12 |
Basic | 97.05 | 94.15 | 93.13 | 93.64 | 88.20 | ||
RandomErasing | 97.87 | 93.96 | 93.95 | 93.96 | 88.61 | ||
Cutout | 97.93 | 94.40 | 93.99 | 94.19 | 89.03 | ||
CutMix | 98.00 | 94.51 | 94.28 | 94.40 | 89.40 | ||
GridMask | 97.92 | 94.81 | 93.51 | 94.16 | 88.96 | ||
Ours | 98.15 | 95.49 | 94.11 | 94.80 | 90.11 | ||
SNE-RoadSeg [fan2020sne] | RGB + Depth | Standard | 95.97 | 92.37 | 91.88 | 92.12 | 86.12 |
Basic | 97.95 | 94.19 | 94.37 | 94.28 | 89.19 | ||
RandomErasing | 98.41 | 95.81 | 95.63 | 95.72 | 91.79 | ||
Cutout | 98.44 | 95.48 | 95.84 | 95.66 | 91.68 | ||
CutMix | 98.30 | 95.16 | 95.86 | 95.51 | 91.40 | ||
GridMask | 98.35 | 94.90 | 95.92 | 95.41 | 91.22 | ||
Ours | 98.66 | 96.89 | 95.54 | 96.21 | 92.70 |
@@ IV-B Dataset
In our experiments, we used the publicly available KITTI road dataset [fritsch2013new] to validate the effectiveness of our algorithm. The KITTI road dataset is one of the most popular and widely used datasets for road scene understanding and is commonly utilized for tasks such as free space detection and lane line detection, it contains 289 frames training data and 290 frames testing data. When using the KITTI road dataset, it is often necessary to divide the dataset into three parts: training dataset, testing dataset, and validation dataset. However, the official KITTI road dataset only provides the training and testing portions and does not include a validation set. Therefore, researchers need to perform their own partitioning of the training and validation sets. Our data division is as follows: First, we randomly split the training data from the KITTI road dataset into two parts, 144 and 145 images, respectively. Out of these, 144 images are used as the testing dataset. Then, from the 145 images, we randomly select 20% of the data, which amounts to 29 images, as the validation dataset, the remaining 116 images are used as the training dataset.
@@ IV-C Experiment Setup
Our experiments are conducted in an Ubuntu 20.04 environment, equipped with an Intel i7 12700F CPU and a NVIDIA GeForce RTX 4090 GPU. We employed the PyTorch framework for model training and set training parameters with a batch size of 2, a total of 300 epochs. Regarding to the augmented data, we used each data augmentation method to generate 3 synthetic images for each origin image in the KITTI road dataset for training. In the standard experimental setup without any data augmentation, we duplicated each original image three times to maintain fairness in the amount of training data.
@@ IV-E Performance Evaluation
The quantitative experimental results on three different architectures of single-modal algorithms U-Net, Swin-UNet, and VM-UNet are shown in Table LABEL:table_single_modal. From the table, it can be seen that our data augmentation method achieves the best performance on three different deep neural network architectures: CNN, Transformer, and Mamba. Specifically, compared to the second best method, on U-Net, our data augmentation method increased the F1-Score from 88.05 to 90.23 and mIoU from 78.65 to 82.20. On Swin-Net, our data augmentation method increased the F1-Score from 86.98 to 87.85 and mIoU from 77.65 to 78.34. On VM-UNet, our data augmentation method improved the F1-Score from 96.03 to 96.22 and mIoU from 92.35 to 92.72. When compared to the Basic method, the F1-Score and mIoU increased by 2.26% and 8.87% on Unet, the F1-Score and mIoU increased by 2.64% and 4.71% on Swin-Unet, and the F1-Score and mIoU increased by 2.05% and 3.94% on VM-Unet.
The quantitative experimental results on two multi-modal algorithms PLARD and SNE-RoadSeg are shown in Table LABEL:table_multi_modal. From the table, it can be seen that our data augmentation method also achieves the best performance on multi-modal models with different data inputs. Specifically, compared to the second best method, on PLARD, our data augmentation method increased the F1-Score from 94.40 to 94.80 and mIoU from 89.40 to 90.11. On SNE-RoadSeg, our data augmentation method increased the F1-Score from 95.72 to 96.21 and mIoU from 91.79 to 92.70. On PLARD, the F1-Score and mIoU increased by 1.24% and 2.17%, on SNE-RoadSeg, the F1-Score and mIoU increased by 2.05% and 3.94%, respectively.
Based on the experiments conducted on both single-modal and multi-modal algorithms mentioned above, our data augmentation methods have shown significant improvements, demonstrating the effectiveness of our approach.

### v2

@@ Abstract
Data augmentation is one of the most common tools in deep learning, underpinning many recent advances including tasks such as classification, detection, and semantic segmentation. The standard approach to data augmentation involves simple transformations like rotation and flipping to generate new images. However, these new images often lack diversity along the main semantic dimensions within the data. Traditional data augmentation methods cannot alter high-level semantic attributes such as the presence of vehicles, trees, and buildings in a scene to enhance data diversity. In recent years, the rapid development of generative models has injected new vitality into the field of data augmentation. In this paper, we address the lack of diversity in data augmentation for road detection task by using a pre-trained text-to-image diffusion model to parameterize image-to-image transformations. Our method involves editing images using these diffusion models to change their semantics. In essence, we achieve this goal by erasing instances of real objects from the original dataset and generating new instances with similar semantics in the erased regions using the diffusion model (as shown in Fig. 1), thereby expanding the original dataset. We evaluate our method on the KITTI road dataset [fritsch2013new] and the Cityscapes dataset [cordts2016cityscapes], and our method achieves the best results compared to other data augmentation methods on both datasets, which demonstrates superiority and effectiveness of our proposed method. Here is our project page: https://sites.google.com/view/data-augmentation.
@@ I INTRODUCTION
In this paper, we propose a novel data augmentation method for free space detection. The method is mainly divided into two steps: The method is mainly divided into two steps: first, using traditional instance segmentation algorithms (such as Mask R-CNN [he2017mask]) or general segmentation algorithms (such as Segment Anything (SAM) [kirillov2023segment]) to erase instance pixels in the background while keeping the foreground pixels unchanged.
Then, the erased regions are locally redrawn using a pre-trained diffusion model to restore the erased parts in the image, as shown in the Fig. 1. During the redrawing process, different linguistic prompts can be used to achieve redrawing of different objects and styles, providing great flexibility. We then test our proposed method on the KITTI road dataset [fritsch2013new] and Cityscapes dataset [cordts2016cityscapes], and the experimental results demonstrate the effectiveness of our approach. Our main contributions are as follows:
@@
We conduct comprehensive experiments on KITTI road dataset and Cityscapes dataset, and the results demonstrate that our data augmentation method achieves the best performance in the free space detection task.
@@ III-B Our Approach
For the “erasing” process, i.e., the masks generation step, we utilized SAM [kirillov2023segment] model, proposed by Menta AI. As a foundational model for image segmentation, SAM demonstrates strong generalization capabilities and performs well across different data domains. We utilize the method proposed in [li2023clip], which involves using text prompts specifying the objects to be erased to generate candidate points on the original image. These candidate points are then used as point prompts for the SAM model to erase objects in the background of the original image. Please note that when erasing background instance pixels, we ensure that foreground pixels remain unaffected, thereby avoiding inconsistencies between the training labels of the generated synthetic data and the original data. As shown in row (f) of Fig. 4, the foreground pixels (free space) of each synthetic data sample remain consistent with the original image. We only erase and redraw elements in the background to alter the original data distribution, thereby achieving the effect of data augmentation.
@@
U-Net [ronneberger2015u] | CNN | Standard | 93.89 | 89.07 | 94.23 | 91.57 | 84.21 |
Basic | 95.24 | 89.35 | 96.90 | 92.97 | 86.87 | ||
RandomErasing | 96.35 | 90.22 | 97.35 | 93.65 | 87.33 | ||
Cutout | 96.59 | 91.21 | 96.88 | 93.96 | 88.54 | ||
CutMix | 96.42 | 90.87 | 96.78 | 93.73 | 88.43 | ||
GridMask | 96.31 | 90.24 | 95.99 | 93.03 | 87.98 | ||
Ours | 97.01 | 92.38 | 97.10 | 94.68 | 89.11 | ||
Swin-UNet [cao2022swin] | Transformer | Standard | 94.31 | 87.25 | 95.17 | 91.03 | 84.22 |
Basic | 95.40 | 89.56 | 97.18 | 93.21 | 87.29 | ||
RandomErasing | 95.88 | 90.35 | 97.98 | 94.01 | 88.34 | ||
Cutout | 96.42 | 91.26 | 97.88 | 94.45 | 88.97 | ||
CutMix | 96.47 | 91.15 | 97.21 | 94.08 | 88.35 | ||
GridMask | 96.13 | 90.98 | 97.35 | 94.06 | 87.98 | ||
Ours | 96.89 | 91.98 | 97.91 | 94.85 | 89.35 | ||
VM-UNet [ruan2024vm] | Mamba | Standard | 95.16 | 93.79 | 95.08 | 94.43 | 89.77 |
Basic | 96.95 | 94.28 | 96.46 | 95.36 | 91.13 | ||
RandomErasing | 97.87 | 96.12 | 97.24 | 96.67 | 92.53 | ||
Cutout | 97.85 | 96.03 | 96.98 | 96.50 | 92.99 | ||
CutMix | 98.15 | 95.78 | 97.33 | 96.55 | 92.87 | ||
GridMask | 97.98 | 96.00 | 97.23 | 96.61 | 92.56 | ||
Ours | 98.55 | 96.77 | 97.51 | 97.14 | 93.15 |
@@ IV-B Dataset
In our experiments, we use the KITTI road dataset [fritsch2013new] and the Cityscapes dataset [cordts2016cityscapes] to validate the effectiveness of our algorithm. The KITTI road dataset is one of the most popular and widely used datasets for road scene understanding, commonly employed for tasks such as free space detection and lane line detection. This dataset comprises 289 frames of training data and 290 frames of testing data. The Cityscapes dataset, on the other hand, focuses on semantic segmentation, instance segmentation, and panoptic segmentation tasks in urban street scenes, encompassing a total of 30 categories. It includes 2975 training images, 500 validation images, and 1525 testing images. Since our work is concentrated on the task of free space detection, we retained only the "road" category from the Cityscapes dataset to validate our method. Since the test labels for both datasets are not public, we partition the original datasets. For the KITTI road dataset, we split the training data into 144 images for testing and 145 images for further division: 20% as validation and the remaining 116 for training. For Cityscapes, we randomly select 50% of the 2975 training images for training and the rest for testing.
@@ IV-C Experiment Setup
Our experiments are conducted in an Ubuntu 20.04 environment, equipped with an Intel i7 12700F CPU and a NVIDIA GeForce RTX 4090 GPU. We employed the PyTorch framework for model training and set training parameters with a batch size of 2, a total of 300 epochs. Regarding to the augmented data, we used each data augmentation method to generate 3 synthetic images for each origin image in datasets for training. In the standard experimental setup without any data augmentation, we duplicated each original image three times to maintain fairness in the amount of training data.
@@ IV-E Performance Evaluation
The quantitative experimental results of three different architectures of single-modal algorithms U-Net, Swin-UNet, and VM-UNet on the KITTI road dataset are shown in Table LABEL:kitti_experiment. From the table, it can be seen that our data augmentation method achieves the best performance on three different deep neural network architectures: CNN, Transformer, and Mamba. Specifically, compared to the second best method, on U-Net, our data augmentation method increased the F1-Score from 88.05 to 90.23 and mIoU from 78.65 to 82.20. On Swin-Net, our data augmentation method increased the F1-Score from 86.98 to 87.85 and mIoU from 77.65 to 78.34. On VM-UNet, our data augmentation method improved the F1-Score from 96.03 to 96.22 and mIoU from 92.35 to 92.72. When compared to the Basic method, the F1-Score and mIoU increased by 2.26% and 8.87% on Unet, the F1-Score and mIoU increased by 2.64% and 4.71% on Swin-Unet, and the F1-Score and mIoU increased by 2.05% and 3.94% on VM-Unet. The quantitative experimental results on Cityscapes dataset [cordts2016cityscapes] are shown in Table III. Similar to the results on the KITTI road dataset, we also achieve the best performance on the Cityscapes dataset.
Based on the experiments conducted on KITTI road dataset and Cityscapes dataset, our data augmentation method has shown promising improvements, demonstrating the effectiveness of our approach.